# HyperEvent: Learning Cohesive Events for Large-scale Dynamic Link Prediction

### v1

@@ Abstract
Dynamic link prediction in continuous-time dynamic graphs is a fundamental task for modeling evolving complex systems. Existing node-centric and event-centric methods focus on individual interactions or atomic states, failing to capture the structural cohesion of composite hyper-events—groups of causally related events. To address this, we propose HyperEvent, a framework reframing dynamic link prediction as hyper-event recognition. Central to HyperEvent is the dynamic construction of an association sequence using event correlation vectors. These vectors quantify pairwise dependencies between the query event and relevant historical events, thereby characterizing the structural cohesion of a potential hyper-event. The framework predicts the occurrence of the query event by evaluating whether it collectively forms a valid hyper-event with these historical events. Notably, HyperEvent outperforms state-of-the-art methods on 4 out of 5 datasets in the official leaderboard. For scalability, we further introduce an efficient parallel training algorithm that segments large event streams to enable concurrent training. Experiments validate HyperEvent’s superior accuracy and efficiency on large-scale graphs. Among which HyperEvent achieves a 6.95% improvement in Mean Reciprocal Rank over state-of-the-art baseline on the large-scale Flight dataset while utilizing only 10.17% of the training time.
Codehttps://github.com/jianjianGJ/HyperEvent
Datasetshttps://tgb.complexdatalab.com/docs/leader˙linkprop
1 Introduction
@@
Temporal graphs model evolving systems through sequences of interactions (i.e., temporal events), serving as essential abstractions for real-world dynamic systems like social networks and transaction networks(Min et al. 2021; Xu et al. 2025; Yan, Weihan, and Chang 2021). These graphs are primarily modeled as discrete-time graph snapshots or Continuous-Time Dynamic Graphs (CTDGs)(Liang et al. 2024). This work focuses specifically on CTDGs because real-world interactions intrinsically manifest as continuous streams of events, where preserving their exact temporal sequence is critical for capturing fine-grained event correlations. Accurately predicting future events within CTDGs, known as dynamic link prediction, remains a core challenge with significant implications for applications such as real-time recommendation(Qin et al. 2024; Zhang et al. 2023), social networks(Daud et al. 2020; Dileo, Zignani, and Gaito 2024), traffic flow prediction(Zou et al. 2024; Ali et al. 2025), knowledge graphs(Ong et al. 2025; Chen et al. 2025), and anomaly detection(Duan et al. 2024; Yang, Zhao, and Shen 2025).
Recent advancements in CTDG representation learning demonstrate significant progress in dynamic link prediction. The evolution of CTDG representation learning has bifurcated into two dominant paradigms: node-centric and event-centric approaches.
Node-centric methods primarily model temporal dynamics through evolving node embeddings. They track node states via RNNs and graph convolutions(Rossi et al. 2020; Kumar, Zhang, and Leskovec 2019; Rakshit Trivedi and Zha 2019; Zhang et al. 2024). While scalable and inductive, these methods often reduce events to pairwise node embeddings, potentially overlooking complex inter-event relationships. Event-centric approaches explicitly model events as first-class entities. Such as CAWN anonymizes temporal motifs via causal walks to capture structural patterns(Wang et al. 2021b). GraphMixer and DyGFormer further reveal that lightweight architectures (e.g., patched Transformers, time-aware MLPs) can effectively encode event sequences without complex RNNs or graph convolutions(Cong et al. 2023; Yu et al. 2023). However, event-centric methods focus on customized embeddings for stacked events and lack explicit mechanisms to model the structural cohesion between events.
Current paradigms overlook a crucial aspect of complex temporal events: the cohesive structure of composite events. Sequences of individual events often have intrinsic causal or synergistic relationships forming larger, meaningful hyper-events. For instance in figure 1, sequential transactions like a credit card purchase at a taxi service ($t_{1}$), followed by food ($t_{2}$), grocery ($t_{3}$), cinema ($t_{4}$), and drink ($t_{5}$) — despite occurring at distinct timestamps — collectively form a cohesive “shopping trip” hyper-event in transaction networks. Existing methods fail to explicitly model these vital event correlations. They treat dynamic link prediction as a function of aggregated but atomized node or node-pair states instead of recognizing the holistic signature of a unified composite hyper-event.
Figure 1: A hyper-event comprising five events.
Moreover, many methods struggle with scalability issues when processing large scale event streams. While node-centric methods offen use a memory-based architecture to store the node embeddings, which require sequential state updates per event, limiting training efficiency and parallelization. Approaches delving deeper into complex event contexts or mutual neighborhood modeling incur even higher computational costs. Consequently, scaling to real-world, large-scale temporal graphs remains a significant challenge.
To overcome these limitations, we propose HyperEvent, an novel framework for dynamic link prediction in CTDGs. HyperEvent fundamentally shifts the predictive task: instead of predicting individual links based solely on node states, it identifies complete hyper-events. A hyper-event is a cohesive set of temporally related events. The core insight is that a query link is likely to occur if and only if it completes a valid hyper-event when combined with relevant historical events. HyperEvent achieves this by introducing an event correlation vector to capture pairwise associations between events. It dynamically constructs a relation sequence characterizing the entire hyper-event without complex parameter learning. This event-level focus naturally captures intrinsic temporal dependencies and boosts prediction accuracy. Furthermore, we develop a highly scalable training algorithm. Leveraging precomputed adjacency tables and segmenting large event streams into parallelizable segments, HyperEvent efficiently trains on real-world, large-scale datasets. This work presents the following contributions:
1) The HyperEvent Framework: HyperEvent presents a novel framework that reframes CTDGs learning for dynamic link prediction, shifting focus from predicting isolated edges to recognizing composite higher-order events.
2) Modeling Event Correlations: HyperEvent introduces an efficiently computed association sequence capturing relational patterns between historical events and the query event. This provides richer contextual information than node-centric or event-centric methods.
3) Scalable Training Algorithm: HyperEvent incorporates an efficient parallel training algorithm, significantly enhancing training efficiency and scalability for large graphs.
4) Empirical Validation: Extensive experiments on multiple benchmark datasets demonstrate HyperEvent’s competitive prediction accuracy and the effectiveness of its parallel training on large-scale graphs.
2 Related Works
Temporal Graph Learning. Research on temporal graph learning has evolved from discrete snapshot aggregation to methods directly processing event streams. Discrete-time techniques partition edge streams into fixed-interval snapshots(Pareja et al. 2020; Manessi, Rozza, and Manzo 2020; Goyal, Chhetri, and Canedo 2020; Hajiramezanali et al. 2019), applying Graph Neural Networks(Scarselli et al. 2008; Kipf and Welling 2017) and temporal models like RNNs(Hochreiter and Schmidhuber 1997; Cho et al. 2014) or Transformers (Vaswani et al. 2017) separately to each snapshot. However, these methods inherently suffer from information loss due to predetermined aggregation granularity, motivating direct processing of continuous event streams.
Node-Centric Continuous-Time Models. Pioneering continuous-time approaches leverage recurrent architectures to maintain node memories. JODIE uses coupled RNNs for mutual user-item updates(Kumar, Zhang, and Leskovec 2019), while DyRep incorporates multi-hop dynamics at increased complexity(Rakshit Trivedi and Zha 2019). The Temporal Graph Network (TGN) framework unifies memory modules with graph propagation(Rossi et al. 2020), adopted by TGAT for temporal attention-based neighbor aggregation(Xu et al. 2020). Despite scalability improvements, these methods remain fundamentally node-centric – sequentially updating node states through event-by-event processing creates computational bottlenecks and constrains parallelization. Recent innovations like TNCN accelerate training via asynchronous propagation(Zhang et al. 2024), and EDGE employs graph partitioning(Chen et al. 2021), yet they retain the core limitation of ignoring holistic event semantics.
Event-Centric Continuous-Time Models. Beyond node embeddings, recent approaches explicitly model event interactions. CAWN encodes causal anonymous walks to capture temporal motifs but incurs prohibitive sampling costs(Wang et al. 2021b). NAT constructs reusable neighbor caches(Luo and Li 2022), while DyGFormer applies Transformers to patched neighbor co-occurrences(Yu et al. 2023). Though these improve topological awareness, they fail to model intrinsic correlations between constituent events within higher-order structures. TCL constructs temporal dependency graphs but inherits RNN limitations for long-range dependencies(Wang et al. 2021a). Continuous-time neural ODE approaches like CTAN enable non-dissipative propagation yet overlook event-level dynamics(Gravina et al. 2024).
3 Method
3.1 Preliminaries and Problem Formulation
@@
Continuous-time Dynamic Graph: A continuous-time dynamic graph represents the dynamics of interactions between entities over a continuous timeline. Formally, it is as a chronologically ordered sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots,(u_{n},v_{n},t_{n})\}$, where $t_{1}\leq t_{2}\leq\cdots\leq t_{n}$. Each event $(u_{i},v_{i},t_{i})$ signifies an interaction occurring at timestamp $t_{i}$ between a source node $u_{i}$ and a destination node $v_{i}$, both belonging to the universal node set $\mathcal{V}$.
@@
Dynamic Link Prediction: The dynamic link prediction task aims to forecast future interactions based on the observed historical evolution of the graph. Given the sequence of all events occurring strictly before a query time $t^{*}$, denoted as $\mathcal{E}_{<t^{*}}=\{(u,v,t)\in\mathcal{G}\mid t<t^{*}\}$, the task is to predict the existence of a specific interaction between nodes $u^{*}$ and $v^{*}$ at time $t^{*}$. Formally, the goal is to determine the likelihood $P\left((u^{*},v^{*},t^{*})\text{ occurs}\mid\mathcal{E}_{<t^{*}}\right)$.
3.2 Existing Frameworks
Figure 2: The HyperEvent framework.
Node-centric methods
@@ 3.2 Existing Frameworks
In practical implementations, the update frequency in equation (1) significantly influences training efficiency, leading to the common approach of updating at intervals of $\Delta t$ as a compromise.
Limitations: This framework faces fundamental constraints preventing parallelization: 1) Causality Lock: Embedding $\mathbf{h}_{u}(t)$ depends strictly on all prior events involving $u$. This sequential dependency forces event-by-event training. 2) Temporal Resolution Dilemma: Ideal granularity ($\Delta t=1$) requires updating node embeddings after every event, maximally preserving recent context but incurring prohibitive $|\mathcal{E}|$ sequential steps for per-epoch training. Practical minibatching sets $\Delta t=\text{batch duration}$, updating embeddings only at batch boundaries. This skips events within the batch interval $[t^{*}-\Delta t,t^{*})$, discarding critical short-term dependencies. Using outdated embeddings $\mathbf{h}(t^{*}-\Delta t)$ to score links at $t^{*}$ ignores crucial events near $t^{*}$ when $\Delta t>1$. This significantly degrades accuracy on high-activity graphs.
Event-centric Methods
@@
where $\phi(\cdot)$ is a heuristic proximity function (e.g., temporal or structural closeness) and $\delta$ a threshold.
The event feature matrix $\mathbf{M}_{e^{*}}\in\mathbb{R}^{|\mathcal{S}_{e^{*}}|\times d}$ is constructed by aggregating features:
@@
Here, $\mathbf{x}$ denotes customized event features, which can be constructed by concatenating edge features and node features, or incorporating other tailored features (such as co-neighbor encoding), along with temporal encoding.
A sequence encoder (e.g., Transformer) processes $\mathbf{M}_{e^{*}}$ to produce the query’s predictive representation:
@@
Limitations: This framework suffers from four critical weaknesses: 1) Feature Dependency: Performance heavily relies on pre-existing node/edge features. When features are absent—common in real-world event streams—manual initialization (e.g., zero vectors) creates informationless inputs, degrading model reliability. 2) Atomic Aggregation: The stacking operation $\mathbf{M}_{e^{*}}$ treats events as isolated entities, disregarding latent inter-event correlations within $\mathcal{S}_{e^{*}}$ that could elucidate complex interaction dynamics. 3) Relational Blindness: By operating solely on raw features, the approach fails to explicitly model structural or temporal dependencies between events, leading to suboptimal context representation. 4) Computational Overhead: The $\phi(\cdot)$-based sampling requires expensive neighborhood searches for every query, incurring $O(|\mathcal{E}|)$ complexity that prohibits scalability to large-scale graphs.
3.3 HyperEvent Framework
HyperEvent fundamentally redefines dynamic link prediction as the verification of hyper-event. Instead of learning stacked isolated event embeddings, it predicts whether a query event $e^{*}=(u^{*},v^{*},t^{*})$ forms a consistent hyper-event $\mathcal{H}_{e^{*}}$ with relevant historical events. A hyper-event represents an integrated interaction pattern where constituent events exhibit intrinsic relational dependencies. The existence probability of $e^{*}$ is derived from the authenticity of $\mathcal{H}_{e^{*}}$:

### v2

# HyperEvent: A Strong Baseline for Dynamic Link Prediction via Relative Structural Encoding

@@ Abstract
Learning representations for continuous-time dynamic graphs is critical for dynamic link prediction. While recent methods have become increasingly complex, the field lacks a strong and informative baseline to reliably gauge progress. This paper proposes HyperEvent, a simple approach that captures relative structural patterns in event sequences through an intuitive encoding mechanism. As a straightforward baseline, HyperEvent leverages relative structural encoding to identify meaningful event sequences without complex parameterization. By combining these interpretable features with a lightweight transformer classifier, HyperEvent reframes link prediction as event structure recognition. Despite its simplicity, HyperEvent achieves competitive results across multiple benchmarks, often matching the performance of more complex models. This work demonstrates that effective modeling can be achieved through simple structural encoding, providing a clear reference point for evaluating future advancements. The code for HyperEvent is accessible on GitHub [^1] https://github.com/jianjianGJ/HyperEvent.
I Introduction
@@
Temporal graphs model evolving systems through sequences of interactions (i.e., temporal events), serving as essential abstractions for real-world dynamic phenomena like social networks and transaction systems[1, 2, 3]. These graphs are primarily modeled as discrete-time graph snapshots or continuous-time dynamic graphs (CTDGs)[4]. This work focuses specifically on CTDGs because real-world interactions intrinsically manifest as continuous streams of events, where preserving their exact temporal sequence is critical for capturing fine-grained event relationships. Accurately predicting future events within CTDGs, known as dynamic link prediction, remains a core challenge with significant implications for applications such as real-time recommendation[5, 6] and anomaly detection[7, 8].
Recent years have witnessed notable advances in representation learning for CTDGs. Proposed methods have aimed to address challenges such as capturing complex structural-temporal dependencies, modeling long-range interactions, and improving generalization to unseen nodes. These approaches can be broadly categorized into two paradigms: node-centric and event-centric methods. Node-centric methods maintain and update memory states for each node, using recurrent or attention-based mechanisms to evolve node embeddings over time[9, 10, 11, 12]. Some incorporate neural ODEs to model continuous dynamics[13], or employ neighbor caching to alleviate structural bottlenecks[14]. While often scalable and inductive, these techniques reduce each event to an update of its endpoint embeddings, potentially overlooking higher-order relationships between events. In contrast, event-centric methods treat interaction events as primary entities. Techniques such as CAWN[15] leverage causal anonymous walks to extract temporal motifs, while DyGFormer[16] uses attention mechanisms to implicitly model event group patterns. Several recent methods also highlight that relatively simple architectures—such as MLPs or patched transformers—can achieve strong performance[17, 16].
Figure 1: An example of a hyper-event (shopping trip) comprising five sequential transaction events (taxi, food, grocery, cinema, drink) along a timeline ($t_{1}$$t_{5}$).
Among these, EdgeBank[18] stands out as a simple, non-parametric baseline that relies on memorizing and replaying historical interactions. However, its performance has been significantly surpassed by newer learning-based models, diminishing its utility as a meaningful benchmark. This creates a critical gap in the literature: the absence of a strong, simple baseline that can serve as a fundamental reference point for evaluating model efficacy. Without such a baseline, it becomes difficult to distinguish genuine advancements from unnecessary complexity, potentially leading to illusory progress.
An effective baseline should be conceptually clear, easy to implement, and competitive with contemporary methods. To this end, this paper introduces HyperEvent, a novel baseline model for dynamic link prediction that combines handcrafted, non-learned feature sequences with a lightweight transformer classifier. The approach is motivated by the observation that individual events often exhibit natural patterns, forming cohesive higher-order structures. For example, as illustrated in Figure 1, a sequence of transactions—such as taking a taxi, purchasing food, buying groceries, watching a movie, and having drinks—may collectively form a natural sequence of related events. HyperEvent leverages this insight by constructing feature sequences that encode relative structural patterns between the query event and relevant historical events. These features serve as input to a standard transformer that classifies whether the query event follows a plausible pattern, and hence is likely to occur.
HyperEvent offers a straightforward approach to event prediction by focusing on relative structural patterns rather than complex event relationships. This simplification allows the model to capture meaningful event sequences without relying on intricate learned embeddings or memory modules. Despite its simplicity, HyperEvent achieves competitive performance compared to state-of-the-art methods across multiple benchmark datasets. It thus fills an important gap as a strong, simple baseline that can aid the community in evaluating model improvements.
The main contributions of this work are summarized as follows:
1) A New Simple Baseline: HyperEvent is proposed as a straightforward baseline for dynamic link prediction that combines interpretable feature construction with a standard classifier, offering a clear reference for future research.
2) Relative Structural Encoding: A simple feature extraction strategy is introduced to capture relative structural patterns between the query event and historical events, enabling the model to recognize meaningful event sequences without complex parameterization.
3) Highly Scalable Inference: The model architecture permits efficient and parallelizable training and inference, making it suitable for large-scale temporal graphs.
4) Extensive Empirical Validation: Comprehensive experiments demonstrate that HyperEvent achieves competitive results across multiple datasets, often matching or exceeding more complex models.
II Related Work
Temporal Graph Learning for Dynamic Link Prediction. Dynamic link prediction is an essential task for forecasting future interactions in continuously evolving networks, with critical applications across social networks [19, 20], communication systems [21, 22], traffic prediction [23, 24], and knowledge graph completion [25, 26]. Unlike traditional static graph methods that rely on structural heuristics or node similarity metrics [27, 28, 29, 30, 31], dynamic settings require modeling temporal event streams, capturing complex dependencies among interactions, and generalizing to unseen nodes [32, 33]. Early approaches discretized event streams into snapshot sequences [34, 35, 36, 37] and applied graph neural networks [38, 39] combined with sequence models such as RNNs [40, 41] or Transformers [42]. However, such discretization often leads to information loss and limited expressiveness, motivating a shift toward continuous-time modeling.
Continuous-Time Dynamic Graph Learning. CTDG methods can be broadly categorized into node-centric and event-centric paradigms. Node-centric approaches maintain and update memory states for each node upon every interaction. Seminal works include JODIE [10], which uses coupled RNNs to update user and item representations, and DyRep [11], which incorporates multi-hop propagation. The Temporal Graph Network (TGN) framework unifies various memory modules and graph aggregation operators[9], and TGAT employs self-attention to aggregate temporal neighbor information[43]. Though later methods such as TNCN [12] and EDGE [44] improve scalability through asynchronous updates or graph partitioning, they remain constrained by sequential, event-by-event processing that limits parallelism and overlooks higher-order event semantics. On the other hand, event-centric methods explicitly model interactions between events. CAWN [15] and TPNet [45] encode anonymous walks to capture temporal motifs. NAT introduces reusable neighbor representations[14], and DyGFormer applies Transformers to patched event sequences[16], implicitly modeling event group patterns through attention mechanisms. TCL constructs temporal dependency graphs but relies on RNNs that struggle with long-range dependencies[46]. Neural ODE-based approaches like CTAN enable continuous dynamics modeling but often ignore fine-grained event correlations[13]. A common trend across these lines of research is the increasing architectural complexity, with many recent models incorporating sophisticated mechanisms such as attention, memory, or neural differential equations. Despite reported improvements, the field lacks a simple, yet effective baseline that clarifies the minimal complexity required for competitive performance, making it difficult to assess whether proposed enhancements genuinely capture essential temporal-topological phenomena or merely introduce inductive biases.
The Role of Baselines and the Limitations of EdgeBank. Strong baselines are fundamental to meaningful scientific progress in machine learning. They serve as essential references that help the community discern actual advancements from false positives, ensure fair comparisons, and maintain a clear lower bound of performance that new methods should exceed. A poorly performing baseline can misleadingly inflate the perceived utility of new models and encourage unnecessary complexity. In temporal graph learning, EdgeBank has been adopted as a canonical baseline—a simple non-parametric method that predicts future edges based on their recurrence in historical event sequences[18]. While computationally efficient and easy to implement, its performance has fallen significantly behind modern learned models across multiple datasets. This growing performance gap diminishes its utility as a meaningful reference, raising concerns about pseudo-progress in the field. When a simple heuristic is substantially outperformed, it becomes unclear whether newly proposed architectures are capturing genuine temporal dynamics or simply leveraging additional parameters and complexity to memorize dataset-specific patterns. Therefore, there is a pressing need for a new, stronger, and principled baseline that is both conceptually simple and highly competitive, enabling more rigorous evaluation and providing clearer insights into what constitutes necessary model complexity for dynamic link prediction.
III Method
III-A Preliminaries and Problem Formulation
@@
Continuous-time Dynamic Graph: A continuous-time dynamic graph represents the dynamics of interactions between entities over a continuous timeline. Formally, it is as a chronologically ordered sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots,(u_{n},v_{n},t_{n})\}$, where $t_{1}\leq t_{2}\leq\cdots\leq t_{n}$. Each event $(u_{i},v_{i},t_{i})$ signifies an interaction occurring at timestamp $t_{i}$ between a source node $u_{i}$ and a destination node $v_{i}$, both belonging to the universal node set $\mathcal{V}$. While events may possess associated features $\mathbf{e}_{u_{i},v_{i}}^{t_{i}}$, our approach primarily focuses on the temporal sequence of interactions for the core problem formulation.
@@
Dynamic Link Prediction: The dynamic link prediction task aims to forecast future interactions based on the observed historical evolution of the graph. Given the sequence of all events occurring strictly before a query time $t^{*}$, denoted as $\mathcal{E}_{<t^{*}}=\{(u,v,t)\in\mathcal{G}\mid t<t^{*}\}$, the task is to predict the existence of a specific interaction between nodes $u^{*}$ and $v^{*}$ at time $t^{*}$. Formally, the goal is to determine the likelihood $P\left((u^{*},v^{*},t^{*})\text{ occurs}\mid\mathcal{E}_{<t^{*}}\right)$. This prediction should be accurate not only for interactions involving nodes observed during the historical period (transductive setting) but also for interactions potentially involving nodes unseen before $t^{*}$ (inductive setting).
III-B Existing Frameworks
III-B1 Node-centric methods
@@
Figure 2: The HyperEvent prediction pipeline: 1)Temporal neighborhood interactions are maintained in real-time adjacency tables; 2) Hyper-events are extracted from source/destination node neighborhoods; 3)Structural-temporal correlations are computed as relational matrices; 4)A discriminator verifies hyper-event authenticity to predict query event occurrence.
III-B2 Event-centric Methods
@@
where $\phi(\cdot)$ is a heuristic proximity function (e.g., temporal or structural closeness) and $\delta$ a threshold.The event feature matrix $\mathbf{M}_{e^{*}}\in\mathbb{R}^{|\mathcal{S}_{e^{*}}|\times d}$ is constructed by aggregating features:
@@
Here, $\mathbf{x}$ denotes customized event features, which can be constructed by concatenating edge features and node features, or incorporating other tailored features (such as co-neighbor encoding), along with temporal encoding.A sequence encoder (e.g., Transformer) processes $\mathbf{M}_{e^{*}}$ to produce the query’s predictive representation:
@@
The Risk of Pseudo-Progress: While these methods have contributed significantly to the field, they often introduce substantial complexity through multi-component architectures involving both spatial and temporal modules. This complexity can obscure whether performance improvements arise from capturing genuine temporal-topological phenomena or simply from increased model capacity. Such drawbacks highlight a risk of pseudo-progress, where increasingly intricate models are proposed without a clear baseline to confirm that the added complexity corresponds to fundamental gains in understanding or capability.
III-C HyperEvent––A Simple Structural Encoding Baseline
In response to the need for a straightforward yet effective baseline in dynamic link prediction, we propose a method based on relative structural encoding. This approach focuses on capturing the structural relationships between a query event and its historical context through carefully designed, handcrafted features.
Intuition for Relative Structural Encoding: Our approach is motivated by the observation that interactions between nodes often follow patterns that can be identified by examining their relative positions within the historical interaction graph. Specifically, we consider the following intuitive notions of structural relationships:1) Direct co-occurrence: How frequently nodes directly interact with the query event’s nodes (0-hop).2) Immediate neighborhood overlap: How the neighborhoods of query event’s nodes overlap with historical events (1-hop).3) Broader neighborhood patterns: How the neighborhoods of neighborhoods overlap, providing a wider context (2-hop).These relative relationships can be computed efficiently and provide a rich representation of the structural context surrounding the query event, which surprisingly proves effective for dynamic link prediction.
Instead of learning stacked isolated event embeddings, HyperEvent predicts whether a query event $e^{*}=(u^{*},v^{*},t^{*})$ forms a consistent hyper-event $\mathcal{H}_{e^{*}}$ with relevant historical events. A hyper-event represents an integrated interaction pattern where constituent events exhibit intrinsic relational dependencies. The existence probability of $e^{*}$ is derived from the authenticity of $\mathcal{H}_{e^{*}}$: